Fill out the Amortization schedule for the following scenario. You set a present value annuity to pay off a $25,000 car in 4 annual installments withan interest of 6%a.Use TVM to compute payments.Fill out the Amortization schedule for the following scenario. You set a present value annuity to pay off a $25,000 car in 4 annual installments withan interest of 6%a.Use TVM to compute payments.
Answer #1

Annual instalment = P(A/P, i, n)

= 25,000(A/P, 6%, 4)

= 25,000(0.2886)

= $7,215